Question – 9
(a) Chlorophyll the green colouring matter of plants contains 2.71% of
magnesium by mass. Calculate the number of magnesium atoms in
3.50 g of chlorophyll. [2.5]
(b) Commercially available aq.Conc.HCl is 38% by mass. Calculate the
molarity of solution.(d=1.19 g/ml). What volume of conc. HCl is
required to make 1L of 0.10 M HCl ? [2.5]
Question – 10
(a) 1.18 g of a mixture of Na2CO3 & K2SO4 is present in 250 ml aqueous
solution. 25 ml of this solution is completely neutralized by 20 ml of
0.05 M H2SO4. Calculate the percentage composition of the
mixture. [2.5]
(b) 4.15 g of an acid are dissolved in water and volume of solution is
made 500 ml by adding water . 20 ml of this solution require 20 ml
of N/10 NaOH solution . If the acid is dibasic, calculate the molar
mass of acid.
1
Expert's answer
2021-08-11T05:18:23-0400

Mass of magnesium in 3.50 g of chlorophyll

= 2.71×3/100

=0.0813g

24.3 g of Mg contain =6.02×1023

atoms

0.0813 g of Mg contain =

24.3

6.02×1023

×0.0813

=1.189×1021atoms

= 1.189×1021/6.02 ×24.3

= 4.8g